# Electronic structure and transport in materials with flat bands: 2D materials and quasicrystals

### v1

@@ 3.3.1 Effective Hamiltonian for defects in a single sublattice of a bipartite lattice
Figure 18: Examples of Total DOS and microscopic conductivity $\sigma_{M}$ in Bernal AB bilayer with selective distribution of defects. (a.1) (a.2) Vacancies randomly distributed on the A1 sublattice. (b.1) (b.2) Vacancies randomly distributed on the B2 sublattice. (c.1) (c.2) Vacancies randomly distributed randomly on B1 and B2 sublattices. In the case $c=5\%$, an isolated midgap states bands is founds. At the border of this bands, i.e., $E\approx t_{1}$, states have a spectacular large conductivity. Calculation with the first-neighbor TB model (see caption of Fig. 14). $c$ is the concentration of defects (vacancies) with respect of the total number of atoms. In panels (a) and (b) the midgap states at $E=0$ (see text) are not included. ${\rm G}_{0}=2e^{2}/h$. Numerical method see A.3. From [354, 160].
@@
In BLG the local environment of A and B atoms is different, indeed the B atoms have only three first neighbors (A atoms) in the same layer, while the A atoms have in addition a neighbor A in the other layer (Fig. 14(a)). Thus the probability that an atom or molecule will stick to an atom A or an atom B should be different, and it is reasonable to think that the functionalization of B atoms is favored. This simple argument has been confirmed by DFT calculations [389] showing that H adsorption energy difference between A site and B site is about $\Delta E=85$ meV in favor of B site, when the number of adsorbates is very low. For a larger number of adsorbates, one can therefore expect competition between two contradictory effects: on the one hand preferential adsorption on the B-sites of the bilayer, and on the other hand adsorption on different sublattices of the same layer as expected in MLG [390, 389]. Indeed in MLG, it exists an interaction between defect states that favors configurations with adsorbates on different sublattices. Such asymmetric adsorption properties between sublattice A and sublattice B have been recently suggested by experimental measurements [386, 388], where the distribution of hydrogen adsorbates on the sublattices is adequately controlled.
@@
Overall, BLG lattice is a bipartite lattice of the two sublattices $\alpha$A1,B2} and $\beta$A2,B1}, from which one expects very specific electronic properties produced by selective functionalization [354, 160, 361]. In a previous section 3.3.1 and Ref. [354], we discussed the limiting cases where adsorbates are randomly distributed only on A1 sublattice (B2 sublattice) of layers 1 (layer 2) while the other layer remains pristine. As shows in Fig.18(a.1) and 18(b.1) its leads to a no gap or a gap, respectively.
@@
On one hand, such a selective functionalization leads to the creation of a gap when sublattice B2 is functionalized. This gap is a fraction of one eV, but of at least 0.5 eV for a concentration $c$ of adsorbates larger than 1% of the total number of atoms. On the other hand, functionalization of sublattice A1 decreases the effective coupling between layers, and thus the conductivity increases when $c$ increases, since the pristine layer is less perturbed by the disordered layer when $c$ increases. These two types of selective functionalization exhibit very different and unusual behaviors. This opens the way to the control of electronic properties through selective functionalization, which is experimentally feasible [386]. However, these extreme cases (A1 or B2 functionalization only) seem too simple to correspond to the experimental sample. Indeed, the complexity of the bipartite BLG lattice requires further theoretical studies of other selective adsorbate distributions. This is why we have also studied a combined functionalization of two sublattices [160]. In particular, we have considered cases where midgap states are coupled to each other (vacancies on A1,B1 sublattices, or vacancies on B1,B2 sublattices) and thus form a midgap band, leading to new diffusivity properties that are not a simple combination of the extreme situations studied in Ref. [354], in which midgap states are not coupled together.
@@
A truly exciting case is the quantum diffusion by the midgap bands formed vacancies randomly distributed in B1 and B2 sublattices [160]. B1- and B2-midgap states are distributed over all the structure with different weights on the atoms A1, A2, B1, and B2 [358]. They form a band since B1-Va midgap states and B2-Va midgap states are coupled by the Hamiltonian (Fig. 18(c.1)). Several regimes are present depending on both energy $E$ and vacancy concentration $c$:
@@
    For small concentrations $c$, e.g., $c=1\%$, there is no gap in the DOS (Fig. 18(c.1)) and states around $E_{D}$ form a narrow midgap states band. The corresponding microscopic conductivity $\sigma_{M}$ presents a plateau at a value independent on $c$, $\sigma_{M}\approx 2\sigma_{M,MLG}$ (Fig. 18(c.2)).
@@
    For high concentrations $c$, e.g., $c=5\%$, the density of states (Fig. 18(c.1)) around $E_{D}$ increases significantly, and as a direct consequence, the plateau of conductivity increases $\sigma_{M}>2\sigma_{M,MLG}$ (Fig. 18(c.2)). As explained above (Sec. 3.3.1), in each layer the gap due to B vacancies (B-Va) increases when $c$ increases, therefore for large $c$ the midgap states bandwidth becomes smaller than the gap, and the midgap states band becomes isolated from other states by small gaps at $|E|\gtrsim t_{1}$, where $t_{1}=0.34$ eV is the interlayer hopping term between first neighbor A atoms. The width of this isolated band is $\Delta w\approx 2t_{1}$, i.e., $E\in[-t_{1},t_{1}]$. For large concentrations $c$, the edge states ($E\approx\pm t_{1}$) have a very exotic conductivity $\sigma_{M}$ which strongly increases when $c$ increases, whereas DOS does not change too much. Roughly speaking this spectacular behavior can be explained by considering the coupling between the B1-Va monolayer midgap states and the B2-Va monolayer midgap states. In monolayer, B-Va midgap states are located on A sublattice around a B vacancy. B-Va midgap states of each layer are not coupled to each other. But, since each A orbital is coupled to an A orbital of the other layer, a B1-Va midgap state is coupled to a B2-Va midgap state, with a hopping term $\gamma_{B_{1}-B_{2}}$. For the smallest $d_{B_{1}-B_{2}}$ distance between the B1-Va and the B2-Va (typically first neighbor $B_{1}$-$B_{2}$) one can estimate that $\gamma_{B_{1}-B_{2}}\approx t_{1}$, and $\gamma_{B_{1}-B_{2}}$ decreases when $d_{B_{1}-B_{2}}$ increases. When $c$ increases, the average $d_{B_{1}-B_{2}}$ distance decreases and thus the average $\gamma_{B_{1}-B_{2}}$ value increases. As a result, by a kind of percolation mechanism between monolayer B-midgap states of the two layers, the conductivity through the BLG increases strongly when $c$ increases.
@@
To conclude, by a systematic studied of selective functionalization [358, 160], we prove theoretically that controlled functionalization can be an excellent way to tune BLG conductivity. This is in agreement with recent experimental results [386, 350, 388] showing that it is possible to control the functionalization with an adsorbate rate of the order of 1% of the total number of atoms and to fabricate single and double side adsorbed bilayer graphene. We find a wide variety of original behaviors and have classified them according to the functionalized sublattices, the adsorbate concentration $c$, and the energy. For example, we give the conditions for opening a mobility gap of several 100 meV. Experimentally, and according to Ref. [386], Hydrogen adsorption on B atoms in one layer is energetically favored. For this reason, the study of the specific cases of B1B2-adsorbates is very interesting. An isolated midgap states band is predicted. Spectacularly, for $c>\sim 1\%$, its edge states have a high electrical conductivity due to the large diffusivity of charge carriers, which deserves further investigation. As the functionalization of atoms can be performed experimentally, one can imagine that those of the B1B2-adsorbates can also be carried out, which makes it possible to control the conductivity.
@@ 3.3.3 Universal conductivity of the midgap states
Figure 19: Midgap states conducity, $\sigma(E=E_{D}=0,\tau_{i})$, as a function of inelastic scattering time $\tau_{i}$ and for different defect concentrations $c$, calculated by the formula (52) of the main text. (a) Vacancies randomly distributed on the atoms B1, (b) Vacancies randomly distributed on the atoms A1. In both cases midgap states are uncoupled states at $E_{D}=0$. In (a) B1 vacancies case these states are isolated by gaps, whereas in (b) A1 Vacancies case they are in the continuum metallic band of the pristine layer (layer 2) [354]. $G_{0}=2e^{2}/h$. Numerical method see A.3 and main text. From [160].
It is also interesting to focus on the conduction by flatband midgap state themselves, i.e., here midgap states at energy $E_{D}=0$ when they are not coupled to each other by the Hamiltonian. This is done, for example, when vacancies are randomly distributed in one of the two sublattices of the MLG [354] or when vacancies are randomly distributed in the $A_{1}$ and $A_{2}$ atoms (or A1 and B2 atoms) in BLG [160].
In these midgap states, the average velocity is zero but conduction is possible due to the inherent quantum fluctuations of the velocity which are due to the interband contributions of the velocity correlation function [391, 159]. Indeed, in the presence of inelastic scattering these fluctuations are modified and do not cancel completely at large times which allows electronic diffusion. It results a non-Boltzmann conductivity, similar to the one found in approximant of quasicrystals (Sec. 2.2) and magic angle twisted bilayer graphene (Sec. 2.3.4), and graphene with particular defects inducing flatbands [391, 159]. The microscopic conductivity $\sigma_{M}$ of those midgap states is negligible for rather small inelastic mean free time $\tau_{i}$. But at large $\tau_{i}$ (large $L_{i}$), the Kubo-Greenwood conductivity of midgap states is [159],
@@
where $n_{i}$ and $\mathcal{D}$ are the DOS and the diffusivity (Eq. (64)) in the presence of inelastic scattering. Since midgap states are non-dispersive states at $E=0$, $n_{i}$ is the broadening of the Delta function, $c\delta(E)$, by a Lorentzian with a width at half maximum $\eta$, $\eta=\hbar/\tau_{i}$. Thus at Dirac energy $E_{D}=0$,
@@
where $S$ is the surface of the unit cell. As an example, the figure 19 shows two extreme cases of conductivity through the mid-gap state at $E=E_{D}$ in the BLG, where resonant defects (vacancies) are randomly distributed in B1 sites or A1 sites, respectively. For the B1 vacancies case (Fig. 19(a)), the midgap states at $E_{D}=0$ are isolated by gaps; and therefore, for large $\tau_{i}$, $\sigma(E=0,\tau_{i})$ reaches a universal constant value, independent of the defect concentration $c$, which equals two times the graphene one [159], $\sigma(E=0)\approx 1.3\,G_{0}$, as it was found also for vacancies in A1, A2 sites, and for vacancies in A1, B2 sites [160]. For A1 vacancies case (Fig. 19(b)), the situation is completely different because the midgap states are located in layer 1 only [354], whereas layer 2 remains pristine. Therefore, at intermediate $\tau_{i}$ values, and for sufficiently large defects concentration $c$, the conductivity of the bilayer is driven by the midgap states plateau values of layer 1, $\sigma(E=0,\tau_{i})\approx 0.65\,G_{0}$. At large $\tau_{i}$ values or small concentrations $c$, the conductivity is dominated by the ballistic conductivity through layer 2 and thus $\sigma(E=0,\tau_{i})\propto\tau_{i}^{2}$. Numerically this contribution of layer 2 may be due to the accumulation of small numerical errors. Yet further investigations are needed as we cannot exclude that this behavior is intrinsic to the system as is observed [159] for the dice model, where the peak of localized states is not in a true gap but just at the edge of the continuum.
@@ 3.4 Metallic carbon nanotubes
Figure 20: 1D / 2D regime transition in metallic carbon nanotubes. Electronic properties metallic single walled carbon nanotube $(60,0)$ with random distribution of resonant defects (vacancies): (a) Density of states (DOS) with different small defect concentration $c=0$, 0.1, 0.2, 0.4% showing a 1D behavior. (b) Localization length $\xi$ versus elastic mean-free path $L_{e}$ showing that Thouless relation is satisfied: $\xi(E_{F})=(N_{\rm ch}(E_{F})+2)L_{e}(E_{F})$ (dashed line), where $N_{\rm ch}(E_{F})$ is close to the number of channels at the energy $E_{F}$ but tends to be higher. Points are for different energies in the corresponding plateau, point colours indicate the concentration $c$ of defects as in panel (a). (c) DOS for $c=1$, 2, 4% showing a 2D behavior, i.e., very similar to (d) monolayer graphene. (e) Ratio $\sigma_{M}/C_{h}$ versus Fermi energy $E_{F}$, for $c=1$%, where $\sigma_{M}$ is the microscopic conductivity and $C_{h}=60a=147,4\,{\rm\AA}$, the circumference of the nanotube, (Sec. 3.1). Depending on $E_{F}$ values the conductivity at large inelastic mean-free path $L_{i}$, $L_{i}\gg L_{e}$, cross over a 1D/2D localization behavior: (f) for small $E_{F}$ value, $\sigma$ is linear in $\ln L_{i}$ (1D law), (g) for large $E_{F}$ value, $\sigma$ is linear in $L_{i}$ (2D law). (f) and (g) are drawn for $c=1$%. The TB model includes only first neighbours with a hopping term $t=2.7$ eV. Dirac energy $E_{D}=0$. $G_{0}={2e^{2}}/{h}$. Numerical method see A.3. From [364, 392].
@@
Numerous studies have been carried out on the electronic transport properties of functionalized carbon nanonubes, including or instance, theoretical [362, 363, 364, 365], and experimental works [393] (and Refs. therein). It is obviously a 1D system, but it exhibits a 1D/2D transition in the presence of a sufficiently large number of defects to be of interest in this review.
@@
The case of metallic nanotubes with resonant defects (therefore with flat bands if the system with defects is periodic) shows results related to graphene. The type of defects, the circumference $C_{h}$ of the nanotube and the Fermi energy are determinant for the localization phenomenon and the transport regime [364]. In short, for a sufficiently low defect concentration $c$, the elastic mean-free path $L_{e}$ of the charge carriers is greater $C_{h}$ and the transport regime is be 1D regime. But when for a sufficiently high $c$, it is possible to have $L_{e}$ and the localization length $\xi$ smaller than the circumference and the regime becomes 2D as for monolayer graphene. This can already be seen in the density of states (DOS), which are very close to the DOS of the free-defect nanotube at low $c$ (Fig. 20(a)) and very similar to monolayer graphene at high $c$ (Fig. 20(c,d)). At low concentrations $c$, transport is well described by the 1D quantum transport theory except very close to the neutrality point ($E_{F}\approx E_{D}$) where even the DOS is modified by the adsorbates. This 1D behavior is confirmed by the linear dependence of $\xi$ on $L_{e}$ (Thouless relation, Fig. 20(b)), which is not achieved at high $c$.
@@
For intermediate concentration $c$ of resonnant defects, this 1D / 2D regime transition exist also when $E_{F}$ varies as illustrated figure 20(e,f,g) for $c=1$%. At low energies, as shown in Fig. 20(f), the scale dependant conductivity follows a decreasing trend (dashed lines), $\sigma_{2D}(L_{i})\approx\sigma_{0}-\alpha G_{0}\ln\left({L_{i}}/{L_{e}}\right)$ (Eq. 34), indicating a logarithmic scale for large $L_{i}$ which is caracteristic of a 2D behaviour [189]. The slope of $\sigma_{2D}(L_{i})$ on this scale is universal, with a value $\alpha\approx{1}/{\pi}$, consistent with observations in 2D monolayer graphene [38]. As shown in Figs. 20(e,g), this 2D behavior occurs when the localization length $\xi$ is smaller than the nanotube circumference $(\xi<C_{h})$. The 1D regime is confirmed by the linear dependence of $\sigma$ on $L_{i}$ for large $L_{i}$, $\sigma(Li)\approx\sigma_{0}-G_{0}(L_{i}-L_{e})$ (dashed lines) [189]. As shown Fig. 20(e), the crossover between 1D and 2D regimes is determined by the ratio ${\xi}/{C_{h}}$. Three distinct regimes can be identified: (1) a 2D regime when $\xi<C_{h}$, (2) an intermediate regime when $\xi\approx C_{h}$, (3) a 1D regime when $\xi>C_{h}$.
@@
The results present Fig. 20 are calculated with TB model including first neighbour only and similar results with nearest-neighbor hopping or even through Variable Range Hopping even at room temperature provided that the Fermi energy is sufficiently close from the charge Neutrality Point and also for sufficiently high concentration of resonant defects [364].

### v2

@@ 3.3.1 Effective Hamiltonian for defects in a single sublattice of a bipartite lattice
Figure 18: Examples of total DOS and microscopic conductivity $\sigma_{M}$ in Bernal AB bilayer with selective distribution of defects. (a.1) (a.2) Vacancies randomly distributed on the A1 sublattice. (b.1) (b.2) Vacancies randomly distributed on the B2 sublattice. (c.1) (c.2) Vacancies randomly distributed randomly on B1 and B2 sublattices. In the case $c=5\%$, an isolated midgap-states band is found. At the border of this bands, i.e., $E\approx t_{1}$, states have a spectacular large conductivity. Calculation with the first-neighbor TB model (see caption of Fig. 14). $c$ is the concentration of defects (vacancies) with respect of the total number of atoms. In panels (a) and (b) the midgap states at $E=0$ (see text) are not included. ${\rm G}_{0}=2e^{2}/h$. For the numerical method see A.3. From Missaoui18 , Jouda21 .
@@
In BLG the local environment of A and B atoms is different, indeed the B atoms have only three first neighbors (A atoms) in the same layer, while the A atoms have in addition a neighbor A in the other layer (Fig. 14(a)). Thus the probability that an atom or molecule will stick to an atom A or an atom B should be different, and it is reasonable to think that the functionalization of B atoms is favored. This simple argument has been confirmed by DFT calculations Moaied14 showing that the H adsorption energy difference between A site and B site is about $\Delta E=85$ meV in favor of B site, when the number of adsorbates is very low. For a larger number of adsorbates, one can therefore expect competition between two contradictory effects: on the one hand preferential adsorption on the B-sites of the bilayer, and on the other hand adsorption on different sublattices of the same layer as expected in MLG Boukhvalov08 , Moaied14 . Indeed in MLG, there exists an interaction between defect states that favors configurations with adsorbates on different sublattices. Such asymmetric adsorption properties between sublattice A and sublattice B have been recently suggested by experimental measurements Katoch18 , Joucken21 , where the distribution of hydrogen adsorbates on the sublattices is adequately controlled.
@@
Overall, the BLG lattice is a bipartite lattice of the two sublattices $\alpha$ {A1,B2} and $\beta$ {A2,B1}, from which one expects very specific electronic properties produced by selective functionalization Missaoui18 , Jouda21 , Kaladzhyan23 . In the previous section 3.3.1 and Ref. Missaoui18 , we discussed the limiting cases where adsorbates are randomly distributed only on A1 sublattice (B2 sublattice) of layers 1 (layer 2) while the other layer remains pristine. As shown in Fig.18(a.1) and 18(b.1) it leads to the absence or the presence of a gap, respectively.
@@
On one hand, such a selective functionalization leads to the creation of a gap when sublattice B2 is functionalized. This gap is a fraction of one eV, but of at least 0.5 eV for a concentration $c$ of adsorbates larger than 1% of the total number of atoms. On the other hand, functionalization of sublattice A1 decreases the effective coupling between layers, and thus the conductivity increases when $c$ increases, since the pristine layer is less perturbed by the disordered layer when $c$ increases. These two types of selective functionalization exhibit very different and unusual behaviors. This opens the way to the control of electronic properties through selective functionalization, which is experimentally feasible Katoch18 . However, these extreme cases (A1 or B2 functionalization only) seem too simple to correspond to the experimental situation. Indeed, the complexity of the bipartite BLG lattice requires further theoretical studies of other selective adsorbate distributions. This is why we have also studied a combined functionalization of two sublattices Jouda21 . In particular, we have considered cases where midgap states are coupled to each other (vacancies on A1, B1 sublattices, or vacancies on B1, B2 sublattices) and thus form a midgap band, leading to new diffusivity properties that are not a simple combination of the extreme situations studied in Ref. Missaoui18 , in which midgap states are not coupled.
@@
A truly exciting case is the quantum diffusion by the midgap bands formed by vacancies randomly distributed in the B1 and B2 sublattices Jouda21 . B1- and B2-midgap states are distributed over the entire structure with different weights on the atoms A1, A2, B1, and B2 Missaoui17 . They form a band since B1-vacancy midgap states and B2-vacancy midgap states are coupled by the Hamiltonian (Fig. 18(c.1)). Several regimes are present depending on both energy $E$ and vacancy concentration $c$:
@@
    For small concentrations $c$, e.g., $c=1\%$, there is no gap in the DOS (Fig. 18(c.1)) and states around $E_{D}$ form a narrow midgap-states band. The corresponding microscopic conductivity $\sigma_{M}$ presents a plateau at a value independent on $c$, $\sigma_{M}\approx 2\sigma_{M,MLG}$ (Fig. 18(c.2)).
@@
    For high concentrations $c$, e.g., $c=5\%$, the density of states (Fig. 18(c.1)) around $E_{D}$ increases significantly, and as a direct consequence, the plateau of conductivity increases $\sigma_{M}>2\sigma_{M,MLG}$ (Fig. 18(c.2)). As explained above (Sec. 3.3.1), in each layer the gap due to B vacancies (B-vacancy) increases when $c$ increases, therefore for large $c$ the midgap-states bandwidth becomes smaller than the gap, and the midgap-states band becomes isolated from other states by small gaps at $|E|\gtrsim t_{1}$, where $t_{1}=0.34$ eV is the interlayer hopping term between first-neighbor A atoms. The width of this isolated band is $\Delta w\approx 2t_{1}$, i.e., $E\in[-t_{1},t_{1}]$. For large concentrations $c$, the edge states ($E\approx\pm t_{1}$) have a very exotic conductivity $\sigma_{M}$ which strongly increases when $c$ increases, whereas the DOS does not change too much. Roughly speaking this spectacular behavior can be explained by considering the coupling between the B1-vacancy monolayer midgap states and the B2-vacancy monolayer midgap states. In a monolayer, B-vacancy midgap states are located on the A sublattice around a B vacancy. B-vacancy midgap states of each layer are not coupled to each other. However, since each A orbital is coupled to an A orbital of the other layer, a B1-vacancy midgap state is coupled to a B2-vacancy midgap state, with a hopping term $\gamma_{B_{1}-B_{2}}$. For the smallest $d_{B_{1}-B_{2}}$ distance between the B1-vacancy and the B2-vacancy (typically first neighbor $B_{1}$-$B_{2}$) one can estimate that $\gamma_{B_{1}-B_{2}}\approx t_{1}$, and $\gamma_{B_{1}-B_{2}}$ decreases when $d_{B_{1}-B_{2}}$ increases. When $c$ increases, the average $d_{B_{1}-B_{2}}$ distance decreases and thus the average $\gamma_{B_{1}-B_{2}}$ value increases. As a result, by a kind of percolation mechanism between monolayer B-midgap states of the two layers, the conductivity through the BLG increases strongly when $c$ increases.
@@
To conclude, by a systematic studied of selective functionalization Missaoui17 , Jouda21 , we prove theoretically that controlled functionalization can be an excellent way to tune BLG conductivity. This is in agreement with recent experimental results Katoch18 , Son20 , Joucken21 showing that it is possible to control the functionalization with an adsorbate concentration of the order of 1% of the total number of atoms and to fabricate single- and double-side adsorbed bilayer graphene. We find a wide variety of original behaviors and have classified them according to the functionalized sublattices, the adsorbate concentration $c$, and the energy. For example, we give the conditions for opening a mobility gap of several 100 meV. Experimentally, and according to Ref. Katoch18 , Hydrogen adsorption on B atoms in one layer is energetically favored. For this reason, the study of the specific cases of B1-B2-adsorbates is very interesting. An isolated midgap-states band is predicted. Spectacularly, for $c\gtrsim 1\%$, its edge states have a high electrical conductivity due to the large diffusivity of charge carriers, which deserves further investigation. As the functionalization of atoms can be performed experimentally, one can imagine that those of the B1-B2-adsorbates can also be carried out, which makes it possible to control the conductivity.
@@ 3.3.3 Universal conductivity of the midgap states
Figure 19: Midgap states conducity, $\sigma(E=E_{D}=0,\tau_{i})$, in Bernal AB bilayer as a function of inelastic scattering time $\tau_{i}$ and for different defect concentrations $c$, calculated by the formula (52) of the main text. (a) Vacancies randomly distributed on the atoms B1, (b) Vacancies randomly distributed on the atoms A1. In both cases midgap states are uncoupled states at $E_{D}=0$. In (a) B1 vacancies case these states are isolated by gaps, whereas in (b) A1 Vacancies case they are in the continuum metallic band of the pristine layer (layer 2) Missaoui18 . $G_{0}=2e^{2}/h$. For the numerical method see A.3 and main text. From Jouda21 .
It is also interesting to focus on the conduction by flat-band midgap state themselves, i.e., here midgap states at energy $E_{D}=0$ when they are not coupled to each other by the Hamiltonian. This is done, for example, when vacancies are randomly distributed in one of the two sublattices of MLG Missaoui18 or when vacancies are randomly distributed in the A1 and A2 atoms (or A1 and B2 atoms) in BLG Jouda21 .
In these midgap states, the average velocity is zero but conduction is possible due to the inherent quantum fluctuations of the velocity which are due to the interband contributions of the velocity correlation function Bouzerar20 , Bouzerar21 . Indeed, in the presence of inelastic scattering these fluctuations are modified and do not cancel completely at large times which allows electronic diffusion. This yields a non-Boltzmann conductivity, similar to the one found in approximants of quasicrystals (Sec. 2.2) and magic-angle twisted bilayer graphene (Sec. 2.3.4), and graphene with particular defects inducing flat bands Bouzerar20 , Bouzerar21 . The microscopic conductivity $\sigma_{M}$ of those midgap states is negligible for rather small inelastic mean free time $\tau_{i}$. However, at large $\tau_{i}$ (large $L_{i}$), the Kubo-Greenwood conductivity of midgap states is Bouzerar21 ,
@@
where $n_{i}$ and $\mathcal{D}$ are the DOS and the diffusivity (Eq. (67)) in the presence of inelastic scattering. Since midgap states are non-dispersive states at $E=0$, $n_{i}$ is the broadening of the Delta function, $c\delta(E)$, by a Lorentzian with a width at half maximum $\eta$, $\eta=\hbar/\tau_{i}$. Thus at the Dirac energy $E_{D}=0$,
@@
where $S$ is the surface of the unit cell. As an example, figure 19 shows two extreme cases of conductivity through the midgap state at $E=E_{D}$ in BLG, where resonant defects (vacancies) are randomly distributed in B1 sites or A1 sites, respectively. For the B1 vacancies case (Fig. 19(a)), the midgap states at $E_{D}=0$ are isolated by gaps; therefore, for large $\tau_{i}$, $\sigma(E=0,\tau_{i})$ reaches a universal constant value, independent of the defect concentration $c$, which equals two times the graphene one Bouzerar21 , $\sigma(E=0)\approx 1.3\,G_{0}$, as it was found also for vacancies in A1, A2 sites, and for vacancies in A1, B2 sites Jouda21 . For the A1 vacancies case (Fig. 19(b)), the situation is completely different because the midgap states are located in layer 1 only Missaoui18 , whereas layer 2 remains pristine. Therefore, at intermediate $\tau_{i}$ values, and for sufficiently large defect concentration $c$, the conductivity of the bilayer is driven by the midgap-states plateau value of layer 1, $\sigma(E=0,\tau_{i})\approx 0.65\,G_{0}$. At large $\tau_{i}$ values or small concentrations $c$, the conductivity is dominated by the ballistic conductivity through layer 2 and thus $\sigma(E=0,\tau_{i})\propto\tau_{i}^{2}$. Numerically this contribution of layer 2 may be due to the accumulation of small numerical errors. Yet further investigations are needed as we cannot exclude that this behavior is intrinsic to the system as is observed Bouzerar21 for the dice model, where the peak of localized states is not in a true gap but just at the edge of the continuum.
@@ 3.4 Metallic carbon nanotubes
Figure 20: 1D / 2D regime transition in metallic carbon nanotubes. Electronic properties of a metallic single-walled carbon nanotube $(60,0)$ with random distribution of resonant defects (vacancies): (a) Density of states (DOS) with different small defect concentration $c=0$, 0.1, 0.2, 0.4% showing a 1D behavior. (b) Localization length $\xi$ versus elastic mean free path $L_{e}$ showing that Thouless relation is satisfied: $\xi(E_{F})=(N_{\rm ch}(E_{F})+2)L_{e}(E_{F})$ (dashed line), where $N_{\rm ch}(E_{F})$ is close to the number of channels at the energy $E_{F}$ even if it tends to be higher. Points are for different energies in the corresponding plateau, point colors indicate the concentration $c$ of defects as in panel (a). (c) DOS for $c=1$, 2, 4% showing a 2D behavior, i.e., very similar to (d) monolayer graphene. (e) Ratio $\sigma_{M}/C_{h}$ versus Fermi energy $E_{F}$, for c=1c=1%, where $\sigma_{M}$ is the microscopic conductivity and $C_{h}=60a=147.4\,{\rm\AA }$, the circumference of the nanotube (Sec. 3.1). Depending on the value of $E_{F}$, the conductivity at large inelastic mean free path $L_{i}$, $L_{i}\gg L_{e}$, crosses over to a 1D/2D localization behavior: (f) for a small value of $E_{F}$, $\sigma$ is linear in $\ln L_{i}$ (2D law), (g) for a large value of $E_{F}$, $\sigma$ is linear in $L_{i}$ (1D law). (f) and (g) are drawn for c=1c=1%. The TB model includes only first neighbors with a hopping term $t=2.7$ eV. Dirac energy $E_{D}=0$. $G_{0}={2e^{2}}/{h}$. For the numerical method see A.3. From Jemai19 , Jemai_these .
@@
Numerous studies have been carried out on the electronic transport properties of functionalized carbon nanotubes, including for instance, theoretical Latil2004 , IshiiInelastic2010 , Jemai19 , Fan20 , and experimental works Jouni22 (and Refs. therein). It is obviously a 1D system, but it exhibits a 1D/2D transition in the presence of a sufficiently large number of defects and thus is of interest in this review.
@@
The case of metallic nanotubes with resonant defects (therefore with flat bands if the system with defects is periodic) shows results related to graphene. The type of defects, the circumference $C_{h}$ of the nanotube and the Fermi energy are determinant for the localization phenomenon and the transport regime Jemai19 . In short, for a sufficiently low defect concentration $c$, the elastic mean free path $L_{e}$ of the charge carriers is greater $C_{h}$ and the transport regime is the 1D regime. However, for a sufficiently high $c$, it is possible to have $L_{e}$ and the localization length $\xi$ smaller than the circumference and the regime becomes 2D as for monolayer graphene. This can already be seen in the density of states (DOS), which is very close to the DOS of the free-defect nanotube at low $c$ (Fig. 20(a)) and very similar to monolayer graphene at high $c$ (Fig. 20(c,d)). At low concentrations $c$, transport is well described by the 1D quantum transport theory except very close to the neutrality point ($E_{F}\approx E_{D}$) where even the DOS is modified by the adsorbates. This 1D behavior is confirmed by the linear dependence of $\xi$ on $L_{e}$ (Thouless relation, Fig. 20(b)), which is not achieved at high $c$.
@@
For intermediate concentration $c$ of resonant defects, this 1D / 2D regime transition exists also when $E_{F}$ varies as illustrated in Fig. 20(e,f,g) for c=1c=1%. At low energies, as shown in Fig. 20(f), the scale-dependent conductivity follows a decreasing trend (dashed lines), $\sigma_{2D}(L_{i})\approx\sigma_{0}-\alpha G_{0}\ln\left({L_{i}}/{L_{e}}\right)$ (Eq. (34)), indicating a logarithmic scale for large $L_{i}$ which is caracteristic of a 2D behavior Lee85 . The slope of $\sigma_{2D}(L_{i})$ on this scale is universal, with a value $\alpha\approx{1}/{\pi}$, consistent with observations in 2D monolayer graphene Trambly13 . As shown in Figs. 20(e,g), this 2D behavior occurs when the localization length $\xi$ is smaller than the nanotube circumference $(\xi<C_{h})$. The 1D regime is confirmed by the linear dependence of $\sigma$ on $L_{i}$ for large $L_{i}$, $\sigma(Li)\approx\sigma_{0}-G_{0}(L_{i}-L_{e})$ (dashed lines) Lee85 . As shown in Fig. 20(e), the crossover between the 1D and 2D regimes is determined by the ratio ${\xi}/{C_{h}}$. Three distinct regimes can be identified: (1) a 2D regime when $\xi<C_{h}$, (2) an intermediate regime when $\xi\approx C_{h}$, (3) a 1D regime when $\xi>C_{h}$.
@@
The results presented in Fig. 20 are calculated with the TB model including first neighbors only and similar results with nearest-neighbor hopping or even variable range hopping even at room temperature, provided that the Fermi energy is sufficiently close from the charge neutrality point and also for sufficiently high concentration of resonant defects Jemai19 .